OpenStudy (gilbert000):

Which value is needed to create a perfect square trinomial from the expression x2 + 8x + _____?

OpenStudy (gunsodo):

x^2+8x+... = x^2+2(4)+... = x^2+2(4)+(...)^2

OpenStudy (gilbert000):

wich one is it

OpenStudy (gunsodo):

You can finish it by yourself. Hint : (x+a)^2 = x^2 + 2ax + a^2

OpenStudy (cecil_the_weasel):

Here is an easy way to do it: Take the second term (the x term), half it, and and then square that. For example: x^2+6x+? 6/2=3 3^2=9 so x^2+6x+9